# Plug-and-Hide: Provable and Adjustable Diffusion Generative Steganography

### 03 Mar 2026 v4

@@ Abstract
Keywords Generative image steganography $\cdot$ Diffusion models $\cdot$ Watermarking
1 Introduction
@@
Steganography, as both an art and a science, aims to conceal secret data within seemingly innocuous carriers such as images for covert communication. In traditional image steganography, benign images (a.k.a. cover images) are usually selected as message carriers, and secret messages are embedded into their least significant bits using either non-adaptive embedding schemes [15] or adaptive embedding strategies [1]. However, such algorithms generally suffer from limited message embedding capacity. Moreover, even at low payloads, message concealment inevitably introduces statistical discrepancies between cover images and their corresponding stego images, increasing the risk of being detected by machine learning-based steganalysis methods [2, 28, 30, 24].
@@
To address these challenges, generative image steganography has emerged as a promising solution. Different from traditional image steganography, generative image steganography employs generative models to directly synthesize stego images from secret messages, thus avoiding the need for pre-defined cover carriers. Early approaches [11, 5, 23, 31, 12, 25, 16, 34, 26, 35] are mainly built upon Generative Adversarial Networks (GANs) [3] and Flow models [9]. GAN-based methods [11, 5, 23, 31, 12, 25, 16, 34] usually require three collaborative networks: a generator, a discriminator, and a message extractor. This demands a delicate trade-off between image generation and message extraction during training, which raises training complexity and often results in unsatisfactory visual quality and message extraction accuracy. In contrast, flow-based methods [26, 35] exploit the strict reversibility of flow models to establish a bijective mapping between secret messages and images. Nevertheless, this inherent reversibility constrains their performance in high-resolution image generation.
@@
Recent studies [29, 6, 8, 33] leverage pre-trained image diffusion models [19, 20, 18] for generative steganography. These methods typically follow a two-step pipeline: the first is an invertible bit-to-noise mapping that transforms secret bits into Gaussian or quasi-Gaussian noise; the second is stego image generation from the noise by solving a probability flow ordinary differential equation (PF-ODE) with solvers such as DPM-Solvers [13, 32, 14]. However, existing diffusion model-based generative image steganography (DM-GIS) methods struggle to flexibly support arbitrary-length secret payloads while simultaneously maintaining high message extraction accuracy, stego image quality, and steganographic security. Specifically, approaches in [29, 6, 33] achieve high extraction reliability and image fidelity but are limited to low effective payloads, which restricts their practicality. Diffusion-Stego [8] supports both low and high payloads but at the expense of compromised security and image quality. These limitations reveal two fundamental issues that remain underexplored: the intrinsic relationship among stego image quality, steganographic security, and message extraction accuracy, as well as the challenge of achieving an optimal trade-off among them. Although Kim et al. empirically identifies a trade-off between image quality and extraction accuracy in DM-GIS [8], it overlooks the critical dimension of steganographic security.
@@
The remainder of this paper is organized as follows. Section˜2 reviews existing DM-GIS algorithms. Section˜3 introduces preliminary knowledge of PF-ODE in diffusion models. Section˜4 analyzes the relationship among stego image quality, steganographic security, and message extraction accuracy, and presents our provable and adjustable PA-B2G mapping. Section˜5 reports the experimental results. Section˜6 concludes this work and discusses its limitations.
2 Related Work
@@
GAN-based Methods. Liu et al. [11] pioneered GAN-based steganography by replacing class labels with secret data to synthesize stego images. Hu et al. [5] employed GANs to convert secret data into noise vectors for generating carrier images. Wang et al. [23] introduced a self-learning GAN-based approach for direct stego image generation from secret messages. Yu et al. [31] proposed an attention-GAN method with optimized network architecture and loss functions. Despite these improvements, such methods still suffer from issues including low image quality and unsatisfactory message extraction accuracy. To address these problems, Liu et al. [12] introduced an image disentanglement autoencoder for generative steganography, which improves both visual quality and extraction accuracy. Wei et al. [25] incorporated a mutual information mechanism to enhance extraction accuracy and designed a hierarchical gradient decay strategy to strengthen anti-steganalysis performance. To tackle robustness problems in message recovery and synthesis quality, Ma et al. [16] developed a weight modulation-based generator and a difference predictor to improve extraction robustness. Most recently, Zhou et al. [34] identified security vulnerabilities in the feature domain of existing GAN-based generative steganography schemes. They further proposed a framework that automatically generates semantic object contours to enhance steganographic security.
@@
Flow-based Methods. To achieve higher embedding capacity and improved message extraction accuracy, Wei et al. [26] proposed a novel flow-based generative steganography algorithm, which directly embeds secret messages into the binarized initial noise of a pre-trained Glow model [9]. In a similar vein, Zhou et al. [35] proposed Secret-to-Image Reversible Transformation (S2IRT), which incorporates the idea of permutation steganography permutation to encode secret bits by rearranging the order of entries in the latent vectors of Glow.
@@
Diffusion Model-based Methods. With the advent of diffusion models, Peng et al. [17] proposed StegaDDPM, a novel DM-GIS algorithm based on the denoising diffusion probabilistic model (DDPM) [4]. StegaDDPM embeds secret information by exploiting the probability distribution between the intermediate states and generated images of DDPM. Yang et al. [29] presented a provably secure bit-to-Gaussian mapping method using distribution-preserving sampling. Hu et al. [6] introduced a novel bit-to-Gaussian mapping module based on orthogonal matrices and dual keys. Zhou et al. [33] proposed a scheme named generative steganography diffusion (GSD) that conceals secret bits in the frequency domain of the initial Gaussian noise. Kim et al. [8] designed four message projection strategies—Message to Noise (MN), Message to Binary (MB), Message to Centered Binary (MC), and Multi-bits—to balance image quality and message extraction performance.
3 Preliminary
@@
Recent studies [21, 4, 7, 22] have demonstrated that visually plausible generated images can be obtained by iteratively denoising pure Gaussian noise. This denoising process can be modeled by the following PF-ODE [22, 7]:
$\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]\mathrm{d}t,$ (1)
@@
where $f(\mathbf{x},t)$ and $g(t)$ denote the drift term and diffusion coefficient with respect to timestep $t$, respectively. $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ denotes the score function [22], which points to the direction of higher probability density of the data $\mathbf{x}$ and is usually approximated by a deep neural network $\frac{1}{\sigma_{t}}\epsilon_{\theta}(\mathbf{x},t)$. Assuming that the integration interval of Equation˜1 is set to $[0,T]$, an image $\mathbf{x}_{0}$ can be generated by integrating Equation˜1 from time $T$ to 0; similarly, $\mathbf{x}_{T}$ can be recovered by integrating Equation˜1 along the reverse time direction. These two computational paths are fully invertible [10]. In practice, the integral associated with the denoising process can be approximately solved using off-the-shelf ODE solvers (e.g., the Euler method [21] and DPM-Solvers [13, 32]) by iteratively denoising pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\sigma_{T}^{2}\mathbf{I})$. This process can be expressed as:
@@
$\mathbf{x}_{t-1}=\mathbf{x}_{t}+\int_{t}^{t-1}\left(f(t)\mathbf{x}_{t}+\frac{g(t)^{2}}{2\sigma_{t}}\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)\right)\,\mathrm{d}t,$ (2)
@@
Likewise, $\mathbf{x}_{T}$ can be approximately recovered by reversing this process. In this paper, we leverage high-order ODE solvers, such as the second-order Heun solver [7] and DPM‑Solver‑2 [13], since they introduce smaller local errors.
4 Our Method
@@
In Section˜4.1, we revisit the fundamental trade-off between stego image quality and message extraction accuracy, originally identified by Kim et al. [8], and further incorporate a discussion of steganographic security. Building upon this analysis, we introduce the PA-B2G mapping in Section˜4.2 and its adjustable variant in Section˜4.3. The detailed message hiding and extraction processes are then presented in Section˜4.4.
4.1 Theoretical Analysis
@@
What Shapes Stego Image Quality and Security? Let $\mathbf{g}_{s}$ and $\mathbf{g}$ denote the noise generated by DM-GIS and pure Gaussian noise, respectively, and let $\mathbf{x}_{s}$ and $\mathbf{x}$ denote the corresponding images generated by the diffusion model with parameter $\theta$. According to Theorem 1 proposed in [36], we have
$D_{\text{KL}}(p(\mathbf{g}_{s})\parallel p(\mathbf{g}))=D_{\text{KL}}(p_{\theta}(\mathbf{x}_{s})\parallel p_{\theta}(\mathbf{x})).$ (3)
@@
The optimization objective of diffusion models can be fundamentally formulated via maximum likelihood estimation [4] and KL divergence, as follows:
$E_{p_{\text{data}}}\left[\log p_{\theta}(\mathbf{x})\right]=D_{\text{KL}}\left(p_{\text{data}}(\mathbf{x})\parallel p_{\theta}(\mathbf{x})\right)+H\left(p_{\text{data}}(\mathbf{x})\right),$ (4)
@@
where $p_{\text{data}}$ and $p_{\theta}$ denote the real image distribution and the learned image distribution, respectively, and $H(\cdot)$ denotes the entropy function. Equation˜4 implies that for a well-trained diffusion model, $D_{\text{KL}}\left(p_{\text{data}}\parallel p_{\theta}\right)=0$. Combined with Equation˜3, we theoretically obtain
$D_{\text{KL}}(p(\mathbf{g}_{s})\parallel p(\mathbf{g}))=D_{\text{KL}}(p_{\theta}(\mathbf{x}_{s})\parallel p_{\text{data}}(\mathbf{x})).$ (5)
@@
Equation˜5 reveals that the Gaussianity of $\mathbf{g}_{s}$ directly affects the stego image quality as well as the steganographic security (see [36] for details). To achieve perfect stego image generation and security, one must ensure that $\mathbf{g}_{s}$ follows a pure Gaussian distribution. Any steganographic operation that compromises the Gaussianity will theoretically degrade both visual quality and security.
@@
What Shapes Message Extraction Accuracy? As analyzed in [36], given a fixed encoding scheme and effective payload, one has to compromise the Gaussianity of $\mathbf{g}_{s}$, such as the adjustment schemes adopted in Diffusion-Stego [8] and S2IRT [35], to achieve higher message extraction accuracy. This inevitably compromises both generation quality and security, as demonstrated in Equation˜5. Notably, to mitigate the destruction of Gaussianity, S2IRT employs an additional perturbation key in its message encoding function, which essentially breaks the constraint of “under the same steganography encoding”.
4.2 Design of PA-B2G
@@
As demonstrated in Section˜4.1, achieving perfect generation quality and security requires that the generated noise $\mathbf{g}_{s}$ be strictly Gaussian noise. To this end, we design a provably invertible bit-to-Gaussian mapping, termed PA-B2G, which comprises two stages. In the first stage, to handle secret bits with arbitrary lengths, the encrypted bit sequence $\mathbf{b}=b_{1}b_{2}\cdots b_{n}$ is divided into several subsequences of length $l$, forming an integer sequence $\mathbf{m}=m_{1}m_{2}\cdots m_{k}$ with $m_{i}\in\{0,1,\cdots,2^{l}-1\}$. Then, the integer sequence $\mathbf{m}$ is transformed into uniformly distributed noise $\mathbf{u}=u_{1}u_{2}\cdots u_{k}$, which can be defined by $\mathcal{C}_{u}(\mathbf{m})=\mathbf{u}$. Inspired by arithmetic coding, we leverage a symmetrical interval partitioning strategy and introduce the following two partition modes:
@@
In the second stage, the obtained sequence $\mathbf{u}$ is transformed into pure Gaussian noise by the inverse transform sampling, which is defined by $\mathbf{g}_{s}=\mathcal{C}_{g}(\mathbf{u})$. $\mathcal{C}_{g}$ is also known as the percent point function (PPF) w.r.t. $\mathcal{N}(0,1)$. The implementation details of PA-B2G and its inverse procedure are provided in Algorithm˜1 and Algorithm˜2, respectively. Note that intervals partitioned by the two modes are all symmetric about $u=\frac{1}{2}$, which is the reason we refer to our method as the symmetrical interval partition, and its rationale is discussed in the next section.
In practice, we can replicate and concatenate the secret message to generate the Gaussian noise $\mathbf{g}_{s}$. This approach essentially leverages repetition coding to improve extraction accuracy at the expense of effective embedding capacity. Consequently, as shown in Section˜5.4, this trade-off is particularly suitable for diffusion model watermarking.
@@
$p(g|m)=p(g|\mathcal{C}_{u}(m))p(\mathcal{C}_{u}(m))=\frac{2^{l}p(g|\mathcal{C}_{u}(m))}{(m\!+\!1\!-\!m)}$ (7)
@@
According to Equation˜6,
$p(g)=p(g|m)p(m)=\frac{1}{2^{l}}\sum\limits_{m}p(g|m)p(m)$ (8)
@@
$p(g|m)=p(g|\mathcal{C}_{u}(m))p(\mathcal{C}_{u}(m))=\frac{2^{l\!+\!1}p(g|\mathcal{C}_{u}(m))}{2(m\!+\!1\!-\!m)}$ (9)
@@
Input: Secret message $\mathbf{m}$.
Output: Pure Gaussian noise $\mathbf{g}$.
1:for $i=1$ to $k$ do
2:  Sample $r_{i}\sim\mathcal{U}(0,1)$;
@@
Input: Noise $\mathbf{g}$.
Output: Secret message $\mathbf{m}$.
1:for $i=1$ to $k$ do
2:  $u_{i}=\mathcal{C}^{-1}_{g}(g_{i})$;
@@
Figure 1: Examples of $l=2$ for modes I and II. Along the vertical axis in each figure, the red regions represent the neighborhoods of quantiles, and the blue regions represent the variance-correction intervals. Sampling within these regions is prohibited. The green regions are the intervals that we actually adjust. Our variance-preserving strategy in each iteration involves two key steps: (a) increasing $S^{2}_{u}$ by increasing $c_{1}$, and (b) decreasing $S^{2}_{u}$ by increasing $c_{2}$.
4.3 Adjustable PA-B2G
Section˜4.2 demonstrates that PA-G2B can generate standard Gaussian noise and achieve provable security. However, in practical scenarios, numerical computation errors in ODE solvers and image quantization both significantly degrade message extraction accuracy. We observe that variations in $u_{i}$ within the $h$-neighborhoods of the quantiles (defined in Theorem˜2) are more likely to induce extraction errors. Therefore, we choose to slightly compromise the Gaussianity of $\mathbf{g}_{s}$ to improve message extraction accuracy.
One straightforward yet effective solution is to avoid sampling within these neighborhoods. According to [8], three conditions must be satisfied to achieve decent image generation performance: i) the sample mean of $\mathbf{g}_{s}$ must be close to 0; ii) the sample variance of $\mathbf{g}_{s}$ must be close to 1; iii) all entries of $\mathbf{g}_{s}$ must be independent and identically distributed. Motivated by these requirements, we propose a variance-preserving algorithm. Before presenting further details, we state the following preliminary results.
Theorem 2.

### 04 Mar 2026 v5

@@ Abstract.
1. Introduction
@@
Steganography, as both an art and a science, aims to conceal secret data within seemingly innocuous carriers such as images for covert communication (Song et al., 2024; Hu et al., 2024a). In traditional image steganography, benign images (a.k.a. cover images) are usually selected as message carriers, and secret messages are embedded into their least significant bits using either non-adaptive embedding schemes (Luo et al., 2010) or adaptive embedding strategies (Filler et al., 2011). However, such algorithms generally suffer from limited message embedding capacity. Moreover, even at low payloads, message concealment inevitably introduces statistical discrepancies between cover images and their corresponding stego images, increasing the risk of being detected by machine learning-based steganalysis methods (Fridrich and Kodovsky, 2012; Xu et al., 2016; You et al., 2020; Wei et al., 2022a).
@@
To address these challenges, generative image steganography has emerged as a promising solution. Different from traditional image steganography, generative image steganography employs generative models to directly synthesize stego images from secret messages, thus avoiding the need for pre-defined cover carriers. Early approaches (Liu et al., 2017; Hu et al., 2018; Wang et al., 2018; Yu et al., 2021; Liu et al., 2022; Wei et al., 2022b; Ma et al., 2023; Zhou et al., 2023; Wei et al., 2022c; Zhou et al., 2022) are mainly built upon Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) and Flow models (Kingma and Dhariwal, 2018). GAN-based methods (Liu et al., 2017; Hu et al., 2018; Wang et al., 2018; Yu et al., 2021; Liu et al., 2022; Wei et al., 2022b; Ma et al., 2023; Zhou et al., 2023) usually require three collaborative networks: a generator, a discriminator, and a message extractor. This demands a delicate trade-off between image generation and message extraction during training, which raises training complexity and often results in unsatisfactory visual quality and message extraction accuracy. In contrast, flow-based methods (Wei et al., 2022c; Zhou et al., 2022) exploit the strict reversibility of flow models to establish a bijective mapping between secret messages and images. Nevertheless, this inherent reversibility constrains their performance in high-resolution image generation.
@@
Recent studies (Yang et al., 2024; Hu et al., 2024b; Kim et al., 2025; Zhou et al., 2025) leverage pre-trained image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022) for generative steganography. These methods typically follow a two-step pipeline: the first is an invertible bit-to-noise mapping that transforms secret bits into Gaussian or quasi-Gaussian noise; the second is stego image generation from the noise by solving a probability flow ordinary differential equation (PF-ODE) with solvers such as DPM-Solvers (Lu et al., 2022; Zheng et al., 2023; Lu et al., 2025). However, existing diffusion model-based generative image steganography (DM-GIS) methods struggle to flexibly support arbitrary-length secret payloads while simultaneously maintaining high message extraction accuracy, stego image quality, and steganographic security. Specifically, approaches in (Yang et al., 2024; Hu et al., 2024b; Zhou et al., 2025) achieve high extraction reliability and image fidelity but are limited to low effective payloads, which restricts their practicality. Diffusion-Stego (Kim et al., 2025) supports both low and high payloads but at the expense of compromised security and image quality. These limitations reveal two fundamental issues that remain underexplored: the intrinsic relationship among stego image quality, steganographic security, and message extraction accuracy, as well as the challenge of achieving an optimal trade-off among them. Although Kim et al. empirically identifies a trade-off between image quality and extraction accuracy in DM-GIS (Kim et al., 2025), it overlooks the critical dimension of steganographic security.
@@
The remainder of this paper is organized as follows. Section 2 reviews existing DM-GIS algorithms. Section 3 introduces preliminary knowledge of PF-ODE in diffusion models. Section 4 analyzes the relationship among stego image quality, steganographic security, and message extraction accuracy, and presents our provable and adjustable PA-B2G mapping. Section 5 reports the experimental results. Section 6 concludes this work and discusses its limitations.
2. Related Work
@@
GAN-based Methods. Liu et al. (Liu et al., 2017) pioneered GAN-based steganography by replacing class labels with secret data to synthesize stego images. Hu et al. (Hu et al., 2018) employed GANs to convert secret data into noise vectors for generating carrier images. Wang et al. (Wang et al., 2018) introduced a self-learning GAN-based approach for direct stego image generation from secret messages. Yu et al. (Yu et al., 2021) proposed an attention-GAN method with optimized network architecture and loss functions. Despite these improvements, such methods still suffer from issues including low image quality and unsatisfactory message extraction accuracy. To address these problems, Liu et al. (Liu et al., 2022) introduced an image disentanglement autoencoder for generative steganography, which improves both visual quality and extraction accuracy. Wei et al. (Wei et al., 2022b) incorporated a mutual information mechanism to enhance extraction accuracy and designed a hierarchical gradient decay strategy to strengthen anti-steganalysis performance. To tackle robustness problems in message recovery and synthesis quality, Ma et al. (Ma et al., 2023) developed a weight modulation-based generator and a difference predictor to improve extraction robustness. Most recently, Zhou et al. (Zhou et al., 2023) identified security vulnerabilities in the feature domain of existing GAN-based generative steganography schemes. They further proposed a framework that automatically generates semantic object contours to enhance steganographic security.
@@
Flow-based Methods. To achieve higher embedding capacity and improved message extraction accuracy, Wei et al. (Wei et al., 2022c) proposed a novel flow-based generative steganography algorithm, which directly embeds secret messages into the binarized initial noise of a pre-trained Glow model (Kingma and Dhariwal, 2018). In a similar vein, Zhou et al. (Zhou et al., 2022) proposed Secret-to-Image Reversible Transformation (S2IRT), which incorporates the idea of permutation steganography permutation to encode secret bits by rearranging the order of entries in the latent vectors of Glow.
@@
Diffusion Model-based Methods. With the advent of diffusion models, Peng et al. (Peng et al., 2023) proposed StegaDDPM, a novel DM-GIS algorithm based on the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). StegaDDPM embeds secret information by exploiting the probability distribution between the intermediate states and generated images of DDPM. Yang et al. (Yang et al., 2024) presented a provably secure bit-to-Gaussian mapping method using distribution-preserving sampling. Hu et al. (Hu et al., 2024b) introduced a novel bit-to-Gaussian mapping module based on orthogonal matrices and dual keys. Zhou et al. (Zhou et al., 2025) proposed a scheme named generative steganography diffusion (GSD) that conceals secret bits in the frequency domain of the initial Gaussian noise. Kim et al. (Kim et al., 2025) designed four message projection strategies—Message to Noise (MN), Message to Binary (MB), Message to Centered Binary (MC), and Multi-bits—to balance image quality and message extraction performance.
3. Preliminary
@@
Recent studies (Song et al., 2020a; Ho et al., 2020; Karras et al., 2022; Song et al., 2020b) have demonstrated that visually plausible generated images can be obtained by iteratively denoising pure Gaussian noise. This denoising process can be modeled by the following PF-ODE (Song et al., 2020b; Karras et al., 2022):
(1) $\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]\mathrm{d}t,$
@@
where $f(\mathbf{x},t)$ and $g(t)$ denote the drift term and diffusion coefficient with respect to timestep $t$, respectively. $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ denotes the score function (Song et al., 2020b), which points to the direction of higher probability density of the data $\mathbf{x}$ and is usually approximated by a deep neural network $\frac{1}{\sigma_{t}}\epsilon_{\theta}(\mathbf{x},t)$. Assuming that the integration interval of Equation 1 is set to $[0,T]$, an image $\mathbf{x}_{0}$ can be generated by integrating Equation 1 from time $T$ to 0; similarly, $\mathbf{x}_{T}$ can be recovered by integrating Equation 1 along the reverse time direction. These two computational paths are fully invertible (Lai et al., 2025). In practice, the integral associated with the denoising process can be approximately solved using off-the-shelf ODE solvers (e.g., the Euler method (Song et al., 2020a) and DPM-Solvers (Lu et al., 2022; Zheng et al., 2023)) by iteratively denoising pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\sigma_{T}^{2}\mathbf{I})$. This process can be expressed as:
@@
(2) $\mathbf{x}_{t-1}=\mathbf{x}_{t}+\int_{t}^{t-1}\left(f(t)\mathbf{x}_{t}+\frac{g(t)^{2}}{2\sigma_{t}}\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)\right)\,\mathrm{d}t,$
@@
Likewise, $\mathbf{x}_{T}$ can be approximately recovered by reversing this process. In this paper, we leverage high-order ODE solvers, such as the second-order Heun solver (Karras et al., 2022) and DPM‑Solver‑2 (Lu et al., 2022), since they introduce smaller local errors.
4. Our Method
@@
In Section 4.1, we revisit the fundamental trade-off between stego image quality and message extraction accuracy, originally identified by Kim et al. (Kim et al., 2025), and further incorporate a discussion of steganographic security. Building upon this analysis, we introduce the PA-B2G mapping in Section 4.2 and its adjustable variant in Section 4.3. The detailed message hiding and extraction processes are then presented in Section 4.4.
4.1. Theoretical Analysis
@@
What Shapes Stego Image Quality and Security? Let $\mathbf{g}_{s}$ and $\mathbf{g}$ denote the noise generated by DM-GIS and pure Gaussian noise, respectively, and let $\mathbf{x}_{s}$ and $\mathbf{x}$ denote the corresponding images generated by the diffusion model with parameter $\theta$. According to Theorem 1 proposed in (Zhu et al., 2026), we have
(3) $D_{\text{KL}}(p(\mathbf{g}_{s})\parallel p(\mathbf{g}))=D_{\text{KL}}(p_{\theta}(\mathbf{x}_{s})\parallel p_{\theta}(\mathbf{x})).$
@@
The optimization objective of diffusion models can be fundamentally formulated via maximum likelihood estimation (Ho et al., 2020) and KL divergence, as follows:
(4) $E_{p_{\text{data}}}\left[\log p_{\theta}(\mathbf{x})\right]=D_{\text{KL}}\left(p_{\text{data}}(\mathbf{x})\parallel p_{\theta}(\mathbf{x})\right)+H\left(p_{\text{data}}(\mathbf{x})\right),$
@@
where $p_{\text{data}}$ and $p_{\theta}$ denote the real image distribution and the learned image distribution, respectively, and $H(\cdot)$ denotes the entropy function. Equation 4 implies that for a well-trained diffusion model, $D_{\text{KL}}\left(p_{\text{data}}\parallel p_{\theta}\right)=0$. Combined with Equation 3, we theoretically obtain
(5) $D_{\text{KL}}(p(\mathbf{g}_{s})\parallel p(\mathbf{g}))=D_{\text{KL}}(p_{\theta}(\mathbf{x}_{s})\parallel p_{\text{data}}(\mathbf{x})).$
@@
Equation 5 reveals that the Gaussianity of $\mathbf{g}_{s}$ directly affects the stego image quality as well as the steganographic security (see (Zhu et al., 2026) for details). To achieve perfect stego image generation and security, one must ensure that $\mathbf{g}_{s}$ follows a pure Gaussian distribution. Any steganographic operation that compromises the Gaussianity will theoretically degrade both visual quality and security.
@@
What Shapes Message Extraction Accuracy? As analyzed in (Zhu et al., 2026), given a fixed encoding scheme and effective payload, one has to compromise the Gaussianity of $\mathbf{g}_{s}$, such as the adjustment schemes adopted in Diffusion-Stego (Kim et al., 2025) and S2IRT (Zhou et al., 2022), to achieve higher message extraction accuracy. This inevitably compromises both generation quality and security, as demonstrated in Equation 5. Notably, to mitigate the destruction of Gaussianity, S2IRT employs an additional perturbation key in its message encoding function, which essentially breaks the constraint of “under the same steganography encoding”.
4.2. Design of PA-B2G
@@
As demonstrated in Section 4.1, achieving perfect generation quality and security requires that the generated noise $\mathbf{g}_{s}$ be strictly Gaussian noise. To this end, we design a provably invertible bit-to-Gaussian mapping, termed PA-B2G, which comprises two stages. In the first stage, to handle secret bits with arbitrary lengths, the encrypted bit sequence $\mathbf{b}=b_{1}b_{2}\cdots b_{n}$ is divided into several subsequences of length $l$, forming an integer sequence $\mathbf{m}=m_{1}m_{2}\cdots m_{k}$ with $m_{i}\in\{0,1,\cdots,2^{l}-1\}$. Then, the integer sequence $\mathbf{m}$ is transformed into uniformly distributed noise $\mathbf{u}=u_{1}u_{2}\cdots u_{k}$, which can be defined by $\mathcal{C}_{u}(\mathbf{m})=\mathbf{u}$. Inspired by arithmetic coding, we leverage a symmetrical interval partitioning strategy and introduce the following two partition modes:
@@
In the second stage, the obtained sequence $\mathbf{u}$ is transformed into pure Gaussian noise by the inverse transform sampling, which is defined by $\mathbf{g}_{s}=\mathcal{C}_{g}(\mathbf{u})$. $\mathcal{C}_{g}$ is also known as the percent point function (PPF) w.r.t. $\mathcal{N}(0,1)$. The implementation details of PA-B2G and its inverse procedure are provided in Algorithm 1 and Algorithm 2, respectively. Note that intervals partitioned by the two modes are all symmetric about $u=\frac{1}{2}$, which is the reason we refer to our method as the symmetrical interval partition, and its rationale is discussed in the next section.
In practice, we can replicate and concatenate the secret message to generate the Gaussian noise $\mathbf{g}_{s}$. This approach essentially leverages repetition coding to improve extraction accuracy at the expense of effective embedding capacity. Consequently, as shown in Section 5.4, this trade-off is particularly suitable for diffusion model watermarking.
@@
(7) $p(g|m)=p(g|\mathcal{C}_{u}(m))p(\mathcal{C}_{u}(m))=\frac{2^{l}p(g|\mathcal{C}_{u}(m))}{(m\!+\!1\!-\!m)}$
@@
According to Equation 6,
(8) $p(g)=p(g|m)p(m)=\frac{1}{2^{l}}\sum\limits_{m}p(g|m)p(m)$
@@
(9) $p(g|m)=p(g|\mathcal{C}_{u}(m))p(\mathcal{C}_{u}(m))=\frac{2^{l\!+\!1}p(g|\mathcal{C}_{u}(m))}{2(m\!+\!1\!-\!m)}$
@@
1:Secret message $\mathbf{m}$.
2:Pure Gaussian noise $\mathbf{g}$.
3:for $i=1$ to $k$ do
4:  Sample $r_{i}\sim\mathcal{U}(0,1)$;
@@
1:Noise $\mathbf{g}$.
2:Secret message $\mathbf{m}$.
3:for $i=1$ to $k$ do
4:  $u_{i}=\mathcal{C}^{-1}_{g}(g_{i})$;
@@
Figure 1. Examples of $l=2$ for modes I and II. Along the vertical axis in each figure, the red regions represent the neighborhoods of quantiles, and the blue regions represent the variance-correction intervals. Sampling within these regions is prohibited. The green regions are the intervals that we actually adjust. Our variance-preserving strategy in each iteration involves two key steps: (a) increasing $S^{2}_{u}$ by increasing $c_{1}$, and (b) decreasing $S^{2}_{u}$ by increasing $c_{2}$.
4.3. Adjustable PA-B2G
Section 4.2 demonstrates that PA-G2B can generate standard Gaussian noise and achieve provable security. However, in practical scenarios, numerical computation errors in ODE solvers and image quantization both significantly degrade message extraction accuracy. We observe that variations in $u_{i}$ within the $h$-neighborhoods of the quantiles (defined in Theorem 4.2) are more likely to induce extraction errors. Therefore, we choose to slightly compromise the Gaussianity of $\mathbf{g}_{s}$ to improve message extraction accuracy.
One straightforward yet effective solution is to avoid sampling within these neighborhoods. According to (Kim et al., 2025), three conditions must be satisfied to achieve decent image generation performance: i) the sample mean of $\mathbf{g}_{s}$ must be close to 0; ii) the sample variance of $\mathbf{g}_{s}$ must be close to 1; iii) all entries of $\mathbf{g}_{s}$ must be independent and identically distributed. Motivated by these requirements, we propose a variance-preserving algorithm. Before presenting further details, we state the following preliminary results.
Theorem 4.2.